The equation x^2+y^2-2xy+2x++1=0 represents a parabola
1
Expert's answer
2020-11-02T18:43:25-0500

"\\
x^{2}+y^{2}-2xy+2x+1=0" is the given equation.


The general equation of 2nd degree in two variables x,y is

"ax^{2}+by^{2}+2hxy+2gx+2fy+c=0"


Comparing this with the given equation we get a=1,b=1,h=-1,g=1,f=0,c=1.


"\Delta=\begin{vmatrix}
 a & h & g\\
 h & b & f\\
g & f & c
\end{vmatrix}" ="\begin{vmatrix}
 1 & -1 & 1 \\
 -1 & 1 & 0\\
1 & 0 & 1
\end{vmatrix}" =1+1*(-1)+1*(-1)=-1"\neq" 0.


"h^{2}-ab=1-1=0" .


Therefore it represents a parabola.